OpenStudy (anonymous):

Complex Variables: Find the Limit using L'Hopital's Rule:

OpenStudy (anonymous):

\[\lim_{z -> i}\frac{ z^9 + z - 2i }{ z^{15} + 1 }\] not sure how to interpret it since it has that -2i.

OpenStudy (loser66):

Woahhh, Quebec friend, your problem is tough, :)

OpenStudy (anonymous):

Did you try using DeMoivre?

OpenStudy (anonymous):

i haven't tried

OpenStudy (anonymous):

Because you should try evaluating it first, for \(z \sim i \)

OpenStudy (anonymous):

\(i = \cos (90°) + i \sin (90°)\)

OpenStudy (anonymous):

actually i figured it out but i'm not sure if this is correct. do l'hopital's once. get:\[\lim_{z->i} \frac{ 9z^8 + 1 }{ 15z^{14} }\] plugging in z = i gives me the textbook's answer

OpenStudy (anonymous):

if it's a correct way to do it**